A triangle has sides of length 5, 7 and 8. A triangle similar to it has a perimeter of 15. What are the length of the sides of this triangle?

Answer by checkley79(3341)

5+7+8=20 is the perimeter of the first triangle.
15/20=3/4 thus the ratio is 3/4 for the similiar triangle.
(3/4)5=15/4=3.75 for the smaller side.
(3/4)7=21/4=5.25 for the next side.
(3/4)8=24/4=6 for the longest side.
Proof:
3.75+5.25+6=15
15=15